The equation of circumcircle of an equilateral triangle is x2+y2+2gx+2fy+c=0 and one vertex of the triangle is (1,1).The equation of incircle of the triangle is


a.4(x2+y2)=g2+f2


b.4(x2+y2)+8gx+8fy=(1-g)(1+3g)+(1-f)(1+3f)


c.4(x2+y2)+8gx+8fy=g2+f2


d.none of these

Grade:12th Pass

1 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Menka,

 

Note that the distance between Centre (-g,-f) nad the Vertex (1,1) = R

In equilateral Triangle r=R/2

 

So Eqn of incircle is (x+g)^2 + (y+f)^2 = r^2 = (R^2)/4 = 1/4 [ (g+1)^2 + (f+1)^2 ]

Simplify this eqn. It is option (B).

 

Best Regards,

Ashwin (IIT Madras).